# Transfer Learning for Molecular Property Predictions from Small Data Sets

### v1

@@ Abstract
Machine learning has emerged as a new tool in chemistry to bypass expensive experiments or quantum-chemical calculations, for example, in high-throughput screening applications. However, many machine learning studies rely on small data sets, making it difficult to efficiently implement powerful deep learning architectures such as message passing neural networks. In this study, we benchmark common machine learning models for the prediction of molecular properties on small data sets, for which the best results are obtained with the message passing neural network PaiNN, as well as SOAP molecular descriptors concatenated to a set of simple molecular descriptors tailored to gradient boosting with regression trees. To further improve the predictive capabilities of PaiNN, we present a transfer learning strategy that uses large data sets to pre-train the respective models and allows to obtain more accurate models after fine-tuning on the original data sets. The pre-training labels are obtained from computationally cheap ab initio or semi-empirical models and corrected by simple linear regression on the target data set to obtain labels that are close to those of the original data. This strategy is tested on the Harvard Oxford Photovoltaics data set (HOPV, HOMO-LUMO-gaps), for which excellent results are obtained, and on the Freesolv data set (solvation energies), where this method is unsuccessful due to a complex underlying learning task and the dissimilar methods used to obtain pre-training and fine-tuning labels. Finally, we find that the final training results do not improve monotonically with the size of the pre-training data set, but pre-training with fewer data points can lead to more biased pre-trained models and higher accuracy after fine-tuning.
keywords:
Machine Learning, Transfer Learning, Message Passing Neural Networks, Semi-Empirical Methods, Density Functional Theory
[…]
{tocentry}
1 Introduction
@@
In the realm of molecular sciences, predicting the properties of molecules stands as a pivotal pursuit across numerous domains, including pharmaceuticals, materials science, and environmental studies. Traditionally, this endeavor has been tackled by quantum chemistry, relying on sophisticated theoretical frameworks and computational simulations to unravel the intricate behaviors of molecules. However, the advent of machine learning (ML) has ushered in a paradigm shift, offering an alternative approach that leverages data-driven methodologies efficiently compute molecular properties.1, 2, 3, 4, 5
@@
One notable application of ML in this domain is its integration into high throughput screening workflows, where it replaces conventional quantum chemistry methods.3 Through the judicious application of ML algorithms, researchers can expedite the process of identifying promising molecules for drug discovery or materials development, thereby accelerating the pace of scientific innovation. However, the powerful deep neural networks (NNs) often used in such approaches usually need large amounts of training data to finally make accurate predictions.
@@
Machine learning algorithms, particularly deep NNs, have exhibited remarkable success in discerning intricate patterns and relationships within data sets of substantial size. In recent developments, the message passing neural network PaiNN has demonstrated highest performance in predicting molecular properties, achieving state-of-the-art results across various datasets.6, 7 Another prevalent approach in ML for molecular sciences entails the utilization of tailored molecular descriptors, such as SOAP,8, 9 in conjunction with traditional ML models like kernel ridge regression (KRR). In general, as a rule of thumb, deep learning architectures are to be preferred when $>$ 103–104 training points are available, and different approaches, such as KRR, should be considered otherwise.10 Yet, in the context of molecular property prediction, data sets often fall short regarding the typically required large amounts of data for deep NNs. This mismatch often presents a formidable obstacle, impeding the efficacy of ML in scenarios characterized by small data sets, which are common in practical applications.11, 12, 13 Bridging this divide necessitates innovative strategies and methodologies tailored to the unique challenges posed by small data sets. In this paper, we present a transfer learning strategy to address this pressing need.
@@
In transfer learning, the knowledge learned in one setting is used to improve performance on a related target task.14 In a first step, known as pre-training, an NN is trained on a large set of data that is closely related to the target task, yielding an NN with pre-optimized parameters. In the second step, termed fine-tuning, the same model is re-trained using data of the target task. Now, since the model is already close to the optimal set of parameters after the pre-training, much less data is needed during fine-tuning to obtain an accurate model for the target task. This strategy has proven useful in various fields, including image classification15 and captioning,16 gaming strategies,17 and social network analysis.18 In the physical sciences, some previous studies have used transfer learning to improve property predictions of molecules and materials.19, 20 Fur such tasks, Sun et al. found that supervised pre-training generally helps, if the pre-training labels are closely aligned with the downstream tasks, but the success of efficient transfer learning strongly depends on the employed data sets, ML models, and their hyperparameters.21 Moreover, a range of studies recently investigated the use of transfer learning to improve atomic force predictions for molecular dynamics simulations.22, 23, 24, 25, 26, 27, 28, 29
@@
In this study, we start by undertaking a rigorous benchmarking, evaluating standard ML techniques in the context of small molecular data sets. In particular, we compare the use of SOAP descriptors tailored to a standard ML algorithm (KRR, NN, or gradient boosting) to the message passing neural network PaiNN for regression. The comparison is done both on the Harvard Organic Photovoltaics (HOPV) data set,30 containing photovoltaic properties of 350 organic molecules, and on the Freesolv data set,31 which contains solvation energies of 643 small organic molecules.
@@
Secondly, we propose a transfer learning strategy as a means of augmenting predictive modeling with small data sets. It involves leveraging computations from computationally cheap ab initio or semi-empirical models coupled with simple linear regression, to generate pre-training labels for PaiNN.
2 Methods
@@
In this study, we use two small common physical chemistry data sets, the Harvard Organic Photovoltaics (HOPV) data set30 and the Freesolv data set.31 The HOPV data set contains 350 molecules and their photovoltaic data obtained from both experiments and quantum-chemical calculations. In this study, we focus on the HOMO-LUMO gaps obtained from PBE0/def2-SVP density functional theory, which is a common ab initio method used for standard chemistry data sets.32, 33, 34, 35, 36, 37, 7 Note that the version of the HOPV data set used by us (HOPV_15_revised_2) contains a single molecule with an Se atom, which we regarded as a structural outlier and therefore removed from the data set prior to training. The Freesolv data set contains experimental and calculated solvation energies of 643 small molecules,31 from which we use the values obtained from ”alchemical” calculations based on molecular dynamics simulations by the Mobley group.38
@@
For benchmarking standard ML methods with HOPV and Freesolv, we employed the PaiNN implementation of the schnetpack Python library,39 as well as the SOAP implementation of the dscribe library40 in junction with either sklearn’s GradientBoostingRegressor (GBoost), kernel ridge regression, or a simple feed-forward NN. The GBoost algorithm profited from an additional set of simple molecular descriptors concatenated to the SOAP descriptor. For NN learning, to limit the size of the NNs, the sparse SOAP descriptors were compressed by principal component analysis, retaining $9$ of the variance.9, 7 The hyperparameters of PaiNN were optimized using an iterative grid search, and the hyperparameters of the SOAP-ML models were optimized using Bayesian optimization. For PaiNN, a learning rate decay and early stopping were employed, and results were averaged over five runs. The final model hyperparameters are listed in the appendix.
@@
In the second part of this study, we present a method to improve the predictions of the message passing neural network PaiNN,6 which was also identified as the best ML model for the data sets at hand (see Results and Discussion), via transfer learning. For this, we need (i) a (small) molecular data set with high-accuracy data of the target property ($y_{\textrm{true}}$), (ii) a large data set of molecular structures to be used for pre-training, and (iii) a computationally inexpensive method from the field of quantum chemistry to compute the target property for the pre-training data. Cheap ab initio methods, like density functional theory in the linear density approximation (LDA-DFT), or semi-empirical methods typically yield results which are qualitatively correct, but quantitatively far off.41 To correct the results obtained from these methods, we use the respective quantum-chemical (QC) method to compute the target property on the original high-quality data set (i), $y_{\textrm{QC}}$, and apply simple linear regression to obtain the coefficients $a$ and $b$ that optimize the equation $(a\cdot y_{\textrm{QC}}+b)-y_{\textrm{true}}=0$ for this data set. Subsequently, we use the same QC method to compute the target property for the structures of the pre-training data set (ii), again $y_{\textrm{QC}}$, and finally obtain the pre-training properties as $y_{\textrm{pt}}=a\cdot y_{\textrm{QC}}+b$, thought to be similar to that of the methods used in (i).
@@
In this study, we use the molecular structures of the OE62 data set as pre-training structures for the HOPV data set, and the molecular structures of the QM9 data set as pre-training structures for Freesolv (ii). The OE62 data set contains crystal structures of 62 k medium-sized organic molecules,37 and the QM9 data set contains DFT-optimized structures of 134 k small molecules.42 To obtain pre-training labels for these data sets (iii), HOMO-LUMO gaps and solvation energies were calculated with LDA-DFT using the ORCA software,43 and tight-binding density functional theory (DFTB) using the GFN2-xTB method44 implemented in the XTB software.45 Solvation energies were obtained as the difference in single point energy of the molecule in vacuum and the molecule with an implicit solvent model for water (conductor-like polarizable continuum model, CPCM in ORCA,46 and generalized Born model with surface area contributions, GBSA in XTB,47, 48, 49 respectively). In Figure 1, we display the scatter plots of the original HOPV and Freesolv data, respectively, against the data obtained from LDA and XTB, respectively, corrected by linear regression. These plots clearly show that the obtained HOMO-LUMO gaps are very close to the original HOPV data, but the solvation energies are quite dissimilar to the original Freesolv data. Furthermore, for both data sets, the properties obtained from LDA are closer to the original data those obtained from XTB. The mean absolute errors of the fits are: 0.05 eV (HOPV LDA), 0.12 eV (HOPV XTB), 1.1 kcal/mol (Freesolv LDA), 1.7 kcal/mol (Freesolv XTB).
@@
Furthermore, we also GBoost trained on HOPV and Freesolv, respectively, to compute HOMO-LUMO gaps for OE62 structures and solvation energies for QM9 structures, respectively, and used these for pre-training PaiNN. The results are then compared to the pre-training described above, in which the pre-training labels are obtained from cheap quantum-chemical methods.
On average, HOMO-LUMO gap computations took approximately 106 s (LDA-DFT), 0.22 s (XTB), and $<$ 0.1 s (GBoost) per data point in OE62, respectively. Solvation energy computations took approximately 70 s (LDA-DFT), 0.15 s (XTB), and $<$ 0.1 s (GBoost) per data point in QM9, respectively, and 1.33 s (XTB) per data point in OE62.
Pre-training on QM9 data was done for 200 epochs (100 k training, 10 k validation, 20 k test data points). Pre-training on OE62 data was done for 250 epochs for LDA and XTB data (50 k training, 5 k validation, 6 k test data points), and for 300 epochs for GBoost data (35 k training, 5 k validation, 5537 testing data points). For pre-training with GBoost data, the pre-training data containing elements which did not occur in the fine-tuning data set were omitted to keep the employed SOAP descriptors reasonably small and therefore minimize memory requirements. Finally, pre-training was done using 35 k training data points (SOAP+SD+GBoost/OE62 for HOPV, and SOAP+SD+GBoost/QM9 for Freesolv) or 30 k data points (SOAP+SD+GBoost/OE62 for Freesolv), respectively. For each method, three models were trained with different random seeds and the best model as evaluated on the test subset was selected for fine-tuning.
3 Results and Discussion
3.1 ML benchmarks
We start by benchmarking standard chemistry ML models for regression on the two small, but high-quality data sets under investigation, HOPV30 and Freesolv,31 listed under (i) in the Methods section. The results are summarized in Table 1 as mean absolute errors (MAEs, root mean squared errors (RMSE) are given in parenthesis). For both data sets, very low errors are obtained with PaiNN, yielding MAEs of 0.20 eV for HOPV and 0.56 kcal/mol for Freesolv, respectively. For HOPV, the SOAP+SD+GBoost model yields even slightly better results with an MAE of 0.19 eV, while 0.86 kcal/mol are obtained for regression on Freesolv. The results obtained with these algorithms are qualitatively similar to those obtained in previous ML studies on HOPV50, 51, 52, 53, 54, 55, 56 and Freesolv.57, 58, 59, 60, 61, 62, 63, 64, 56, 65, 66, 67

### v2

@@ Abstract
Machine learning has emerged as a new tool in chemistry to bypass expensive experiments or quantum-chemical calculations, for example, in high-throughput screening applications. However, many machine learning studies rely on small data sets, making it difficult to efficiently implement powerful deep learning architectures such as message passing neural networks. In this study, we benchmark common machine learning models for the prediction of molecular properties on two small data sets, for which the best results are obtained with the message passing neural network PaiNN, as well as SOAP molecular descriptors concatenated to a set of simple molecular descriptors tailored to gradient boosting with regression trees. To further improve the predictive capabilities of PaiNN, we present a transfer learning strategy that uses large data sets to pre-train the respective models and allows to obtain more accurate models after fine-tuning on the original data sets. The pre-training labels are obtained from computationally cheap ab initio or semi-empirical models and
both data sets are normalized to mean zero and standard deviation one to align the labels’ distributions. This study covers two small chemistry data sets, the Harvard Organic Photovoltaics data set (HOPV, HOMO-LUMO-gaps), for which excellent results are obtained, and on the Freesolv data set (solvation energies), where this method is less successful, probably due to a complex underlying learning task and the dissimilar methods used to obtain pre-training and fine-tuning labels. Finally, we find that for the HOPV data set, the final training results do not improve monotonically with the size of the pre-training data set, but pre-training with fewer data points can lead to more biased pre-trained models and higher accuracy after fine-tuning.
preprint: AIP/123-QED
I Introduction
@@
In the realm of molecular sciences, predicting the properties of molecules stands as a pivotal pursuit across numerous domains, including pharmaceuticals, materials science, and environmental studies. Traditionally, this endeavor has been tackled by quantum chemistry, relying on sophisticated theoretical frameworks and computational simulations to unravel the intricate behaviors of molecules. However, the advent of machine learning (ML) has ushered in a paradigm shift, offering an alternative approach that leverages data-driven methodologies efficiently compute molecular properties.Montavon et al. (2013); Hansen et al. (2015); Dral (2020); Fung et al. (2021); Singh et al. (2022)
@@
One notable application of ML in this domain is its integration into high throughput screening workflows, where it replaces conventional quantum chemistry methods.Dral (2020) Through the judicious application of ML algorithms, researchers can expedite the process of identifying promising molecules for drug discovery or materials development, thereby accelerating the pace of scientific innovation. However, the powerful deep neural networks (NNs) often used in such approaches usually need large amounts of training data to finally make accurate predictions.
@@
Machine learning algorithms, particularly deep NNs, have exhibited remarkable success in discerning intricate patterns and relationships within data sets of substantial size. In recent developments, the message passing neural network PaiNN has demonstrated highest performance in predicting molecular properties, achieving state-of-the-art results across various datasets.Schütt, Unke, and Gastegger (2021); Kirschbaum et al. (2023) Another prevalent approach in ML for molecular sciences entails the utilization of tailored molecular descriptors, such as SOAP,Bartók, Kondor, and Csányi (2013); Parsaeifard et al. (2021) in conjunction with traditional ML models like kernel ridge regression (KRR). In general, as a rule of thumb, deep learning architectures are to be preferred when $>$ 103–104 training points are available, and different approaches, such as KRR, should be considered otherwise.Unke et al. (2021) Accordingly, molecular fingerprinting methods are commonly used in junction with KRR for regression on small data sets.Zhao et al. (2020, 2022); Kirschbaum et al. (2023); O. Dral (2024) Yet, in the context of molecular property prediction, data sets often fall short regarding the typically required large amounts of data for deep NNs. This mismatch often presents a formidable obstacle, impeding the efficacy of ML in scenarios characterized by small data sets, which are common in practical applications.Ramakrishnan et al. (2015); Zhang and Ling (2018); Mazouin, Schöpfer, and Lilienfeld (2022) Bridging this divide necessitates innovative strategies and methodologies tailored to the unique challenges posed by small data sets. In this paper, we present a transfer learning strategy to address this pressing need.
@@
In transfer learning, the knowledge learned in one setting is used to improve performance on a related target task.Zhuang et al. (2020) In a first step, known as pre-training, an NN is trained on a large set of data that is closely related to the target task, yielding an NN with pre-optimized parameters. In the second step, termed fine-tuning, the same model is re-trained using data of the target task. Now, since the model is already close to the optimal set of parameters after the pre-training, much less data is needed during fine-tuning to obtain an accurate model for the target task. This strategy has proven useful in various fields, including image classificationSaenko et al. (2010) and captioning,Vinyals et al. (2015) gaming strategies,Sharma et al. (2007) and social network analysis.Tang et al. (2016) In the physical sciences, a range of studies recently investigated the use of transfer learning to improve atomic force predictions for molecular dynamics simulations.Smith et al. (2019); Käser et al. (2020); Zheng et al. (2021); Käser, Richardson, and Meuwly (2022); Gao et al. (2022); Zhang et al. (2022); Chen et al. (2023); Zaverkin et al. (2023) Moreover, some previous studies have used transfer learning to improve property predictions of molecules and materials.Li et al. (2022a); Shui et al. (2022) For example, Su et al. recently presented a transfer learning strategy for predicting frontier orbital energies of organic materials based on DFT data obtained using two different functionals for generating pre-training and fine-tuning labels, respectively.Su et al. (2023) Furthermore, several studies have explored the use of transfer learning for regression on both HOPVZhao et al. (2020); Eibeck et al. (2021); Munshi et al. (2021); Li et al. (2022b); Moore, Bardagot, and Banerji (2022) and Freesolv.Vermeire and Green (2021); Li et al. (2022b); Zhang, Xia, and Zhang (2022) For such tasks, Sun et al. found that supervised pre-training generally helps, if the pre-training labels are closely aligned with the downstream tasks, but the success of efficient transfer learning strongly depends on the employed data sets, ML models, and their hyperparameters.Sun, Dai, and Yu (2022) Hoffmann et al. further showed that increasing both the size of the pre-training and the fine-tuning data sets linearly improved the prediction results in the transfer learning settings investigated by them.Hoffmann et al. (2023) However, as pointed out by Dral, relatively little research is done in exploring the advantages and limitations of transfer learning in the context of repurposing the models for different levels and molecules.O. Dral (2024)
@@
In this study, we start by undertaking a rigorous benchmarking, evaluating standard ML techniques in the context of small molecular data sets. In particular, we compare the use of SOAP descriptors tailored to a standard ML algorithm (KRR, NN, or gradient boosting) to the message passing neural network PaiNN for regression. The comparison is done both on the Harvard Organic Photovoltaics (HOPV) data set,Lopez et al. (2016) containing photovoltaic properties of 350 organic molecules, and on the Freesolv data set,Duarte Ramos Matos et al. (2017) which contains solvation energies of 643 small organic molecules.
@@
Secondly, we propose a transfer learning strategy as a means of augmenting predictive modeling with small data sets. It involves leveraging computations from computationally cheap ab initio or semi-empirical models to generate pre-training labels for PaiNN.
II Methods
@@
In this study, we use two small common physical chemistry data sets, the Harvard Organic Photovoltaics (HOPV) data setLopez et al. (2016) and the Freesolv data set.Duarte Ramos Matos et al. (2017) The HOPV data set contains 350 molecules and their photovoltaic data obtained from both experiments and quantum-chemical calculations. In this study, we focus on the HOMO-LUMO gaps obtained from PBE0/def2-SVP density functional theory, which is a common ab initio method used for standard chemistry data sets.Adamo and Barone (1999); Perdew, Ernzerhof, and Burke (1996); Schäfer, Horn, and Ahlrichs (1992); Weigend and Ahlrichs (2005); Rupp et al. (2012); Stuke et al. (2020); Kirschbaum et al. (2023) Note that the version of the HOPV data set used by us (HOPV_15_revised_2) contains a single molecule with an Se atom, which we regarded as a structural outlier and therefore removed from the data set prior to training. The Freesolv data set contains experimental and calculated solvation energies of 643 small molecules,Duarte Ramos Matos et al. (2017) from which we use the values obtained from "alchemical" calculations based on molecular dynamics simulations by the Mobley group.Mobley and Guthrie (2014)
For benchmarking standard ML methods with HOPV and Freesolv, we employed the PaiNN implementation of the schnetpack Python library.Schütt et al. (2019) PaiNN was recently proposed as an extension of the popular MPNN SchNet.Schütt, Unke, and Gastegger (2021) In the PaiNN architecture, molecular atoms are represented as graph nodes, and interactions between atoms are represented as undirected edges between the nodes. Each atom (node) is described by atomic scalar features $s$, and each interaction (edge) as edge vector features $h$ that are updated using stacked message-passing and update blocks. After the message phasing phase, the final predictions are obtained after summing or averaging over the final node descriptors.Schütt, Unke, and Gastegger (2021)
@@
We further used the SOAP implementation of the dscribe libraryHimanen et al. (2020) in junction with either sklearn’s GradientBoostingRegressor (GBoost), kernel ridge regression, or a simple feed-forward NN. The GBoost algorithm profited from an additional set of simple molecular descriptors concatenated to the SOAP descriptor. For NN learning, to limit the size of the NNs, the sparse SOAP descriptors were compressed by principal component analysis, retaining $9$ of the variance.Parsaeifard et al. (2021); Kirschbaum et al. (2023) The hyperparameters of PaiNN were optimized using an iterative grid search, and the hyperparameters of the SOAP-ML models were optimized using Bayesian optimization. For PaiNN, a learning rate decay and early stopping were employed, and results were averaged over five runs. In all cases, 20 % of the data was used for testing, and for the NNs, 20 % of the data was used for validation and early stopping. The final model hyperparameters are listed in the appendix.
@@
In the second part of this study, we present a method to improve the predictions of the message passing neural network PaiNN,Schütt, Unke, and Gastegger (2021) which was also identified as the best ML model for the data sets at hand (see Results and Discussion), via transfer learning. For this, we need (i) a (small) molecular data set with high-accuracy data of the target property ($y_{\textrm{true}}$), (ii) a large data set of molecular structures to be used for pre-training, and (iii) a computationally inexpensive method from the field of quantum chemistry to compute the target property for the pre-training data. Cheap ab initio methods, like density functional theory in the linear density approximation (LDA-DFT), or semi-empirical methods typically yield results which are qualitatively correct, but quantitatively far off.Thiel (2014) To correct for the biases inherent in the results obtained from these methods, we normalize both data sets to have mean 0 and standard deviation 1 before NN training.
@@
In this study, we use the molecular structures of the OE62 data set as pre-training structures for the HOPV data set, and the molecular structures of the QM9 data set as pre-training structures for Freesolv (ii). The OE62 data set contains crystal structures of 62 k medium-sized organic molecules,Stuke et al. (2020) and the QM9 data set contains DFT-optimized structures of 134 k small molecules.Ramakrishnan et al. (2014) To obtain pre-training labels for these data sets (iii), HOMO-LUMO gaps and solvation energies were calculated with LDA-DFT using the ORCA software,Neese et al. (2020) and tight-binding density functional theory (DFTB) using the GFN2-xTB methodBannwarth, Ehlert, and Grimme (2019) implemented in the XTB software.Bannwarth et al. (2021) Solvation energies were obtained as the difference in single point energy of the molecule in vacuum and the molecule with an implicit solvent model for water (conductor-like polarizable continuum model, CPCM in ORCA,Barone and Cossi (1998) and generalized Born model with surface area contributions, GBSA in XTB,Onufriev, Bashford, and Case (2004); Sigalov, Fenley, and Onufriev (2006); Lange and Herbert (2012) respectively). Note that we do not use the properties originally published with the respective data sets QM9 and OE62 during pre-training. To illustrate the data quality obtained from LDA and XTB computations, we display the scatter plots of the original HOPV and Freesolv data, respectively, against the data obtained from LDA and XTB, respectively, corrected by linear regression (Fig. 1). These plots clearly show that the obtained HOMO-LUMO gaps are very close to the original HOPV data, but the solvation energies are quite dissimilar to the original Freesolv data. Furthermore, for both data sets, the properties obtained from LDA are closer to the original data those obtained from XTB. The mean absolute errors of the fits are: 0.05 eV (HOPV LDA), 0.12 eV (HOPV XTB), 1.1 kcal/mol (Freesolv LDA), 1.7 kcal/mol (Freesolv XTB).
@@
On average, HOMO-LUMO gap computations took approximately 106 s (LDA-DFT), and 0.22 s (XTB) per data point in OE62, respectively. Solvation energy computations took approximately 70 s (LDA-DFT), and 0.15 s (XTB) per data point in QM9, respectively, and 1.33 s (XTB) per data point in OE62.
Pre-training on QM9 data was done for 200 epochs (100 k training, 10 k validation, 20 k test data points). Pre-training on OE62 data was done for 250 epochs for LDA and XTB data (50 k training, 5 k validation, 6 k test data points). For each method, three models were trained with different random seeds and the best model as evaluated on the test subset was selected for fine-tuning.
III Results and Discussion
III.1 ML benchmarks
We start by benchmarking standard chemistry ML models for regression on the two small, but high-quality data sets under investigation, HOPVLopez et al. (2016) and Freesolv,Duarte Ramos Matos et al. (2017) listed under (i) in the Methods section. The results are summarized in Table 1 as mean absolute errors (MAEs, root mean squared errors (RMSE) are given in parenthesis). For both data sets, very low errors are obtained with PaiNN, yielding MAEs of 0.01 eV for HOPV and 0.56 kcal/mol for Freesolv, respectively. For HOPV, the SOAP+SD+GBoost model yields an MAE of 0.19 eV, while 0.86 kcal/mol are obtained for regression on Freesolv. The results obtained with these algorithms are qualitatively similar to those obtained in previous ML studies on HOPVPaul et al. (2019); Zhao et al. (2020); Munshi et al. (2021); Eibeck et al. (2021); Fare et al. (2022); Moore, Bardagot, and Banerji (2022); Li et al. (2022b) and Freesolv.Hutchinson and Kobayashi (2019); Scheen et al. (2020); Weinreich, Browning, and von Lilienfeld (2021); Borgis et al. (2021); Vermeire and Green (2021); Gao et al. (2021); Pathak, Mehta, and Priyakumar (2021); Chung et al. (2022); Li et al. (2022b); Low, Coote, and Izgorodina (2022); Zhang, Xia, and Zhang (2022); Zhang et al. (2023) The remaining models SOAP+NN and SOAP+KRR perform considerably worse on both data sets, in direct comparison SOAP+NN yielding better results for HOPV while SOAP+KRR yields better results for Freesolv.